GRADUATE STUDENT COURSE ASSIGNMENTS

Block 1 Discrete mathematics —- Babai  Weeks 1–8
Block 2a Random walks —- Lawler  Weeks 1–2
Block 2b Fourier series —- Fefferman  Weeks 3–4
Block 2c Music and mathematics —- Fiore  Weeks 5–6
Block 3a Mapping class groups —- Farb  Weeks 1–4
Block 3b 2x2 matrices —- Bader  Weeks 5–6
Block 3c Toric geometry —- Kerr  Weeks 7–8
Block 4a TQFT’s —- May  Weeks 1–4
Block 4b Products on spaces —- Angeltveit  Weeks 5–8
Block 6 Number Theory (Apprentice) —- Babai Weeks 1*–4
Week 1 Week 2 Week 3 Week 4 Week 5 Week 6 Week 7 Week 8
Block 1 Block 1 Block 1 Block 1  Block 1 Block 1 Block 1
Block 2a Block 2a Block 2b Block 2b Block 2c Block 2c
Block 3a Block 3a Block 3a Block 3a Block 3b Block 3b Block 3c Block 3c
Block 4a Block 4a Block 4a Block 4a Block 4b Block 4b Block 4b Block 4b
*  Block 5 Block 5 Block 5

*No class meetings first week, but meetings with students

Block 1 (Babai) (weeks 1–4, 6–8)

Balaji, Drenning, Dzhafarov, Marcus, Schedler Block 2a (Lawler) (weeks 1–2)

Beceanu, Peng Block 2b (Fefferman) (weeks 3–4)

Beceanu, Peng Block 2c (Fiore) (weeks 5–6)

Malestein, Thiyagarajan Block 3a (Farb) (weeks 1–4)

Bou-Rabee, Church, Day, Malestein, Putman Block 3b (Bader) (weeks 5–6)

Anders, Dowdall Block 3c (Kerr) (weeks 7–8)

Anders, Peng Block 4a (May) (weeks 1–4)

Bohmann, Boyarchenko, Guillou, Johnson, Lind Block 4b (Angeltveit) (weeks 5–8)

Bohmann, Johnson, Lind Block 5 (Apprentice) (weeks 2–4)

Broshi, Geline, Kerr, Morris, Shipman

2 GRADUATE STUDENT COURSE ASSIGNMENTS

Assignments by block and by mentoring group

Name (28)  weeks  Block  Mentoring
Alan Anders  8  3b, 3c  Apprentice
Sundeep Balaji  8  1  YSP
Marius Beceanu  2-5  2a, 2b  SESAME
Ian Biringer  1-2, 8?  YSP
Michael Broshi  1-6, 8  5  YSP
Anna-Marie Bohmann  8  4  SESAME
Khalid Bou-Rabee  8  3a  Apprentice
Mitya Boyarchenko  8  4a  YSP
Tom Church  8  3a  SESAME
Matthew Day  8  3a  SESAME
Spencer Dowdall  3-8  3a, 3b  YSP
Shawn Drenning  1-7  1  SESAME
Damir Dzhalil Dzhafarov  2-8  1  SESAME
Jim Fowler  8  3a  YSP
Michael Geline  1-4  5  Apprentice
Bertrand Guillou  1-4  4a  Apprentice
Shawn Henry  SESAME, YSP
Niles Johnson  8  4  SESAME
Gabriel Kerr  1-6  5  Apprentice
John Lind  8  4  YSP
Justin Malestein  8  2c 3a  YSP
Aaron Marcus  8  1  YSP (+ 1 sit-in)
Courtney Morris  2-4 or 6  5  Apprentice
Irine Peng  8  2a, 2b, 3c  SESAME
Andrew Putman  1-4  3a  Apprentice
Travis Schedler  2-8  1  SESAME
Ian Shipman  8  5  SESAME (+ 1 sit-in)
Saravanan Thiyagarajan  8  2c  Apprentice

ORGANIZERS and COORDINATORS

Team leaders for courses: people in charge of organizing problem sessions, making sure that graduate students are at talks, and coordinating with faculty, whose preferences on how graduate students are to help out are paramount.

1 Babai: Schedler 2a Lawler: Peng 2b Fefferman: Peng 2c Fiore: Malestein 3a Farb: Putman 3b Bader: Anders 3c Kerr: Anders 4a May: Guillou 4b Angeltveit: Johnson 5 Babai (Apprentice): Broshi

Overall mentoring coordinator: Boyarchenko

Apprentice mentoring: Coordinator: Geline Mentors: Anders, Bou-Rabee, Geline, Guillou, Kerr, Morris, Putman, Thiyagarjan (Reorganization for weeks 5–8 depending on which undergraduates participate)

SESAME mentoring: Coordinator: Johnson Mentors: Beceanu, Bohmann, Church, Day, Drenning, Dzhafarov, Henry, Johnson, Peng, Schedler, Shipman

YSP mentoring: Coordinator: Fowler Mentors: Balaji, Biringer, Broshi, Boyarchenko, Dowdall, Fowler, Henry, Lind, Malestein, Marcus